### v1

# The Critique of Critique

Abstract
Critique, as a natural language description for assessing the quality of model-generated content, has been proven to play an essential role in the training, evaluation, and refinement of Large Language Models (LLMs). However, there is a lack of principled understanding in evaluating the quality of the critique itself. In this paper, we pioneer the critique of critique, termed MetaCritique, which is a framework to evaluate the critique from two aspects, i.e., factuality as precision score and comprehensiveness as recall score. We calculate the harmonic mean of precision and recall as the overall rating called F1 score. To obtain a reliable evaluation outcome, we propose Atomic Information Units (AIUs), which describe the critique in a more fine-grained manner. MetaCritique takes each AIU into account and aggregates each AIU’s judgment for the overall score. Moreover, given the evaluation process involves intricate reasoning, our MetaCritique provides a natural language rationale to support each judgment. We construct a meta-evaluation dataset containing 300 critiques (2653 AIUs) across four tasks (question answering, reasoning, entailment, and summarization), and we conduct a comparative study to demonstrate the feasibility and effectiveness. Experiments also show superior critique judged by MetaCritique leads to better refinement, indicating generative artificial intelligence indeed has the potential to be significantly advanced with our MetaCritique. We will release relevant code and meta-evaluation datasets at https://github.com/GAIR-NLP/MetaCritique.
1 Introduction
Natural language critique has assumed a crucial role in advancing the development of Large Language Models (LLMs), ranging from the training of a more helpful and harmless model (Bai et al., 2022; OpenAI, 2023; Scheurer et al., 2023; Wu et al., 2023), alignment evaluation of model generations (Wang et al., 2023b; Zheng et al., 2023; Chan et al., 2023; Li et al., 2023) to the refinement of defective model outputs (Madaan et al., 2023; Gou et al., 2023; Ye et al., 2023; Akyurek et al., 2023). While a bunch of recent works are being done using generated critiques to assist in the development of large language models (Cui et al., 2023; Kim et al., 2023; Wang et al., 2023a; Li et al., 2023; Ke et al., 2023), there has not been enough emphasis on how to automatically and efficiently evaluate the quality of these critiques due to the following challenges: (i) quantification: establishing specific criteria to determine the critique rating, (ii) reliability: ensuring transparency to calculate the comparable score, and (iii) intricacy: grasping the complex relations among multiple concepts in the critique evaluation.
In this paper, we pioneer the critique of critique, termed MetaCritique, to get over the above hurdles. An example of our MetaCritique is shown in Figure 1. Firstly, our MetaCritique tackles the quantification issue by establishing the criteria for critique evaluation, i.e., a meaningful critique should provide factual statements and comprehensive assessments, which will be quantified by two metrics we developed: precision and recall. Precision serves to gauge the accuracy of the critique’s content, ensuring each point is factual, while recall measures the extent to which the critique fully covers the necessary breadth of information, reflecting its comprehensiveness. Secondly, our MetaCritique addresses the reliability concern by introducing Atomic Information Units (AIUs). AIUs symbolize the fundamental segments of informative critique that can not be divided further. Our MetaCritique converts the critique-level evaluation into AIU-level evaluation to minimize ambiguity in the evaluation process. Subsequently, MetaCritique aggregates these AIU-level outcomes to produce the overall score, thereby guaranteeing the transparency of the scoring process. Lastly, inspired by how complex reasoning problems are alleviated using LLMs (Zhou et al., 2023), we attempt to resolve the intricacy problem by generating a natural language rationale step-by-step, i.e., Chain-of-Thought (CoT) (Wei et al., 2022), for each AIU-level judgment. As a result, it enhances the reliability of judgment and facilitates human involvement in the evaluation loop.
To validate the feasibility and effectiveness of our MetaCritique, we create a meta-evaluation dataset covering four tasks (question answering, reasoning, entailment, and summarization) across 16 public datasets, and we conduct a comparative analysis of human-written and LLM-generated critiques. Our MetaCritique achieved near-human performance so MetaCritique can serve as an effective critique evaluation for human annotation and LLM’s reflexion. Besides, MetaCritique can identify high-quality critiques, which lead to improved results via iterative refinement. It indicates that MetaCritique can enhance the alignment of existing LLMs.
Our contributions are summarized as follows:
We are the pioneers in prioritizing critique evaluation and introducing the critique of critique, termed MetaCritique, involving the principles of precision and recall.
Our MetaCritique is more interpretable and transparent, wherein critiques are decomposed into AIUs, and natural language rationale is provided for each judgment.
We collect a meta-evaluation dataset involving question answering, reasoning, entailment, and summarization to confirm the feasibility and effectiveness of our MetaCritique, which have been publicly released.
2 MetaCritique
MetaCritique aims to evaluate the quality of hypothesis critique by generating the critique of hypothesis critique. It involves three steps, i.e., reference generation, AIU extraction, and hypothesis critiquing as illustrated in Figure 2. To facilitate the detailed description of MetaCritique, we first introduce some important concepts.
Question denotes a user’s query or instruction that prompts the generative model to produce a pertinent and insightful response.
Model-Generated Answer describes the textual content created by the generative model as a reaction to the question.
Reference Answer is a ground-truth answer for the question. It is used for the precision score.
Hypothesis Critique is the natural language feedback to point out errors of the model-generated answer and provide actionable suggestions. It can be written by either human annotators or LLMs. It will be evaluated in this paper.
Reference Critique represents an ideal critique to comment accurately and thoroughly on the model-generated answers. It is used to calculate the recall score.
Atomic Information Unit (AIU) is the smallest unit that can self-sufficiently convey a piece of information without any additional context. MetaCritique decomposes a critique into AIUs to reduce the ambiguity of the evaluation process and improve the transparency of the outcome. Generating a numeric score for the critique depends on the information volume it encompasses. For example, a precision score considers the volume of correct information in the critique. However, information, being intangible, renders the task of determining the volume of information unfeasible (Porat, 1977; Soofi, 1994). By extracting and counting AIUs from a critique, we can approximate the volume of information as the number of AIUs because AIUs are the most fundamental elements.
2.1 Reference Generation
Reference is essential for most text generation evaluation (Papineni et al., 2002; Lin, 2004; Zhang* et al., 2020; Yuan et al., 2021). Also, it is crucial for our MetaCritique. We need reference answers to calculate the precision score, and reference critiques to calculate the recall score. However, reference answers or reference critiques are always unavailable because they require significant human effort. To address this issue, we adopt the GPT-4 generated content as a proxy of the references as in previous works (OpenAI, 2023; Zheng et al., 2023; Peng et al., 2023; Cui et al., 2023; Li et al., 2023).
Considering that GPT-4 can make mistakes with improper instructions, we refined our instruction via manual review, whereby we tested various instructions on a series of examples and assessed the alignment between the generations with human preference. Given appropriate prompts, GPT-4 can achieve remarkable results (6.51/7 for reference answer and 6.56/7 for reference critique under the human evaluation in Table 3). We provide detailed prompting instructions in Table 7.
2.2 AIU Extraction
AIU extraction aims to split a critique into AIUs. It is similar to some prior works such as verifiable claims extraction for factuality detection (Chern et al., 2023) and atomic content units extraction for summarization evaluation (Liu et al., 2023b). It has been proven that suitably prompted GPT-4 can precisely conduct such a task as mentioned in Chern et al. (2023). Inspired by their success, we implement AIUs extraction by prompting GPT-4. We provide detailed instructions in Table 9.
2.3 Hypothesis Critiquing
Precision
We devise precision to verify the factuality of the hypothesis critique. It is motivated by the fact that good critiques should state factual information without any hallucination. Specifically, we design the precision task, which is a binary classification task on the AIU level to validate whether each AIU is factual or non-factual. This task receives the question, model-generated answer and the optional reference answer as the context, and outputs whether an AIU from the hypothesis critique is factual or non-factual along with a natural language rationale.
We prompt strong LLMs with strong instruction-following capability to implement this evaluation. We follow the idea of CoT reasoning to design the instructions for prompting LLMs. Firstly, the LLM needs to find the necessary information to verify the AIU from the question, model-generated answer, and reference answer. Then, the LLM attempts to explain and reason about whether the AIU is factual or not. Finally, the LLM states the conclusion. Detailed prompting instructions with demonstrations are in Table 10.
After checking each AIU in the hypothesis critique, we aggregate the precision score as the proportion of factual AIUs relative to the total count of AIUs in the hypothesis critique as follows:
where is the number of all AIUs extracted from the hypothesis critique, and is the -th AIU of the hypothesis critique.
Recall
We use recall to assess the coverage of the hypothesis critique over the reference critique. It is motivated by the fact that good critiques should contain all key points of the reference critique without any omissions. In this evaluation, we design recall task, which is a binary classification task on the AIU level to classify whether the hypothesis critique entails each AIU of the reference critique. This task receives the hypothesis critique as the premise and outputs whether an AIU from the reference critique is entailed in the hypothesis critique or not along with a natural language rationale.
We prompt strong LLMs with strong instruction-following capability to implement this evaluation. We follow the idea of CoT reasoning to design the instructions for prompting LLMs. Firstly, the LLM attempts to explain and reason about whether the AIU from the reference critique is mentioned or logically inferred from the hypothesis critique. Then, the LLM states the conclusion whether the AIU is entailed or not. Detailed prompting instructions with demonstrations are in Table 11.
After checking each AIU in the reference critique, we denote the recall score of the hypothesis critique as the ratio of entailed AIUs to all AIUs from the reference critique as follows:
where is the number of all AIUs extracted from the reference critique and is the -th AIU of the reference critique.
F1 Score
We introduce the F1 Score as an overall assessment score of hypothesis critique to harmonize the precision and recall. It is calculated as the harmonic mean of the precision score and recall score as follows:
3 Meta-Evaluation Dataset
In this section, we elaborate on constructing a meta-evaluation dataset for comparing MetaCritique with its counterparts, along with presenting its statistical features.
3.1 Collection of Question and Model-Generated Answer
To get broad coverage of NLP domains, we collect ready-made question-answer pairs from the Shepherd (Wang et al., 2023a) dataset, which consists of question-answer-critique triads for generating critique across various domains. The model-generated answer in this dataset has reasonable errors, so that the critique can be generated to improve the response. We carefully extract some data as our meta-evaluation data. As a result, we collect data from four domains: entailment, reasoning, question answering, and summarization, across 16 datasets: Entailment Bank (Dalvi et al., 2021), e-SNLI (Camburu et al., 2018), Adversarial NLI (Nie et al., 2020), ECQA (Aggarwal et al., 2021), CosmosQA (Huang et al., 2019), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), PIQA (Bisk et al., 2020), ARC (Clark et al., 2018), RACE (Lai et al., 2017), SIQA (Sap et al., 2019), TriviaQA (Joshi et al., 2017), Natural Question (Kwiatkowski et al., 2019), BoolQ (Clark et al., 2019), GPT-3 summarization (Goyal et al., 2022), DeFacto (Liu et al., 2023a).
3.2 Collection of Hypothesis Critique
Hypothesis critique is the evaluation object in this paper. It can be written by human annotators (abbreviated to Hypo.h in tables) and LLMs (abbreviated to Hypo.l in tables). It is valuable to evaluate human-written critique, as it can help to understand the common shortcomings of human annotations. Also, it is necessary to understand how well the LLM-generated critiques perform because a large number of LLM-generated critiques have been used for training, evaluation, and refinement.
For human-written critique, we collect the ready-made human-written critique from Shepherd dataset. There is only one human-written critique for a question-answer pair. For LLM-generated critiques, we collect two critiques for each question-answer pair. These two critiques will be further used to conduct pairwise comparisons. We do not conduct the pairwise comparison between human-written critiques and LLM-generated critiques to avoid potentially misleading or unfair outcomes that could arise from inherent differences in their writing styles. In detail, we randomly select two different LLMs from WizardLM (Xu et al., 2023) (13B and 70B), GPT-3.5, LLaMA-2 (Touvron et al., 2023) (chat-13B and chat-70B), Vicuna (Chiang et al., 2023) (13B and 33B), and SelFee (Ye et al., 2023) (13B). Then, we respectively prompt the selected LLM to generate the critique. We use the same prompt as the reference critique generation (Table 7) for GPT-3.5. As for other LLMs, detailed prompting instructions can be found in Table 8.
3.3 Collection of Human Annotation
We collect human labeling results for precision and recall tasks driven by two objectives: (1) to evaluate the performance of different LLMs to execute the precision and recall tasks. (2) to meta-evaluate MetaCritique and its counterparts in the light of human judgments.
Before annotation, we prepare the reference answer, reference critique and extracted AIUs via GPT-4 as introduced in Section 2.1 and 2.2. We engaged two postgraduate students for performing the precision and recall task. As illustrated in Figure 2, they replace the GPT-4 to provide solely binary labels without explanation. A third postgraduate meticulously reviews the work of the first two annotators, ensuring accuracy and resolving any discrepancies. This structured process is essential for maintaining the reliability of our research data. Lastly, we can calculate the precision score, recall score, and F1 score via these annotated results. These scores, representing human judgments, are named as gold scores in this paper.
| Type | # Pair | # Critique | Avg. # AIUs |
| Hypo.h | 100 | 100 | 3.31 |
| Hypo.l | 100 | 200 | 8.10 |
| Reference | 100 | 100 | 7.02 |
3.4 Statistics
Table 1 shows the statistics of our meta-evaluation dataset. The number of AIUs can be viewed as the information volume of the critique. We find that human-written critiques contain less information than LLM-generated critiques (including reference critiques), nearly half of LLM-generated critiques.
| AIU Level (Micro) | |||
| Meta-P | Meta-R | Meta-F1 | |
| Hypo.h | 87.61 | 48.72 | 62.62 |
| Hypo.l | 71.85 | 53.28 | 61.19 |
| Critique Level (Macro) | |||
| Meta-P | Meta-R | Meta-F1 | |
| Hypo.h | 85.37 | 50.97 | 58.24 |
| Hypo.l | 71.07 | 54.37 | 58.20 |
According to human annotation, we calculate the gold scores of precision, recall, and F1 score at the AIU level and critique level. The results are shown in Table 2. We find that human-written critiques get higher precision scores yet lower recall scores than LLM-generated critiques. It indicates that LLMs are even more prone to creating fictional information, while human is hardly possible to make such mistakes. Also, it is worth noting that human misses more key information than LLMs. This might be because human-written critiques produce fewer AIUs than LLM-generated critiques.
4 Experiments
In this section, we conduct five experiments to demonstrate the feasibility and effectiveness of MetaCritique. The Exp. I shows the feasibility of using GPT-4 to generate reference and extract AIUs via human evaluation. The Exp. II shows the capability of various LLMs to execute the precision and recall task. The Exp. III-V show the superiority of MetaCritique variants against two GPT-4 based methods introduced by Wang et al. (2023a).
4.1 Baseline
We compare multiple modern LLMs with strong instruction-following capabilities for AIU-level precision and recall task (i.e., Exp. II). The tested models include WizardLM, LLaMA-2 Chat, Vicuna, GPT 3.5, and GPT-4. Moreover, we randomly choose 100 AIUs and engage a postgraduate student (not the annotator) to perform the same task. This result can serve as an upper benchmark (named as Human). We also introduce a baseline named GPT-4 w/o ans, where we generate the reference answer and perform the precision task in one step. It aims to confirm the importance of pre-generating a reference answer.
We compare four variants of MetaCritique with two GPT-4 based methods for the Exp. III-V:
Meta-P is the precision score of MetaCritique.
Meta-R is the recall score of MetaCritique.
Meta-F1 is the F1 score of MetaCritique.
Meta-Open is the F1 score of MetaCritique powered by open-source LLMs, whereby WizardLM 70B is used for the precision task, and WizardLM 13B is used for the recall task because they beat other open-source LLMs in the Table 4.
Pairwise is to compare two hypothesis critiques via GPT-4 and pick up the better one. The prompting instruction is shown in Table 15.
Single is to generate a likert score (1-7) for a hypothesis critique via GPT-4. The prompting instruction is shown in Table 16.
| Reference Generation | ||
| Answer | Critique | |
| 6.51 | 6.56 | |
| AIUs Extraction | ||
| Hypo.h | Hypo.l | Reference |
| 6.72 | 6.57 | 6.79 |
4.2 Exp-I: Human Evaluation for Reference Generation and AIUs Extraction
Setup
We conduct a human evaluation to validate the quality of the reference answer, reference critique, and extracted AIUs that were generated by GPT-4. We ask each human annotator to rate the outcome on a 1–7 likert score. Detailed instructions for human to evaluate reference answer generation, reference critique generation, and AIUs extraction can be found in Table 12, 13, and 14, respectively. For reference quality annotation, we present as much information as possible to annotators, including the question, model-generated answer, and hypothesis critiques. Also, we allow annotators to carefully search online whenever they need help. As for AIUs quality annotation, we only provide the original critique.
Results
Table 3 shows the human evaluation for the quality of GPT-4 generated answers, critiques, and AIUs. The upper half results that GPT-4 attains remarkable performance can confirm the feasibility of using GPT-4 generated outcomes as references. Moreover, the lower half results that GPT-4 delivers impressive results can justify its use to extract AIUs for MetaCritique. 111We highly suggest using GPT-4 for reference generation and AIUs extraction for MetaCritique.
| Precision Task | Recall Task | ||||
| Model | Size | Hypo.h | Hypo.l | Hypo.h | Hypo.l |
| Human | – | 90.00 | 86.00 | 86.00 | 85.00 |
| Llama 2 Chat | 7B | 31.42 | 34.14 | 8.55 | 5.98 |
| 13B | 56.50 | 52.22 | 55.13 | 51.78 | |
| 70B | 54.08 | 58.09 | 72.08 | 65.46 | |
| Vicuna | 7B | 71.00 | 70.37 | 59.69 | 60.83 |
| 13B | 72.81 | 69.14 | 71.94 | 70.66 | |
| 33B | 62.84 | 60.12 | 72.93 | 63.82 | |
| WizardLM | 7B | 26.89 | 30.19 | 52.28 | 55.56 |
| 13B | 62.54 | 59.81 | 74.50 | 74.22 | |
| 30B | 56.50 | 63.89 | 63.53 | 61.68 | |
| 70B | 79.76 | 72.28 | 72.51 | 71.44 | |
| GPT-3.5 | – | 81.87 | 77.28 | 80.63 | 81.41 |
| GPT-4 w/o ans | – | 86.40 | 81.48 | – | – |
| GPT-4 | – | 89.12 | 87.96 | 85.47 | 86.82 |
| Pearson | Spearman | Kendall’s Tau | ||||
| Hypo.h | Hypo.l | Hypo.h | Hypo.l | Hypo.h | Hypo.l | |
| Single | 0.508 | 0.390 | 0.503 | 0.379 | 0.396 | 0.290 |
| Meta-Open | 0.730 | 0.709 | 0.745 | 0.690 | 0.566 | 0.506 |
| Meta-P | 0.355 | 0.667 | 0.283 | 0.681 | 0.227 | 0.504 |
| Meta-R | 0.844 | 0.831 | 0.837 | 0.830 | 0.681 | 0.649 |
| Meta-F1 | 0.841 | 0.886 | 0.836 | 0.899 | 0.675 | 0.724 |
4.3 Exp-II: AIU-level Accuracy
Setup
Our MetaCritique centers around two binary classification tasks: precision and recall. In this experiment, we investigated the capability of various LLMs to execute two tasks. Each AIU was treated as a unique test case.
Results
We present the AIU-level accuracy results in Table 4. We find that GPT-4 outperforms all models by a large margin and achieves performance (nearly 90%) comparable to that of human, so it is reasonable to use MetaCritique powered by GPT-4 for automated evaluation of human-written critiques and LLM-generated critiques. 222It is highly suggested to implement MetaCritique via GPT-4 due to its impressive performance in aiu-level accuracy. Moreover, WizardLM stands out as the top-performing open-source model, closely rivaling GPT-3.5. Lastly, the degradation (around 3% and 6%) without the reference answer indicates that it is necessary to pre-generate a reference answer.
4.4 Exp-III: Correlation Coefficient
Setup
In this experiment, we use different methods to generate a score for hypothesis critique. We calculate correlation coefficients to measure the correlation between different scoring baselines and human judgments. Specifically, we use Pearson correlation (Lee Rodgers and Nicewander, 1988; Mukaka, 2012), Spearman correlation (Zar, 2005), and Kendall’s Tau (Kendall, 1938) as metrics. We calculate the above correlation coefficients between the outcome score and the gold F1 score. To perform a rigorous analysis, we adopt the bootstrapping method (Koehn, 2004) for significance tests.
Results
In Table 5, we show the correlation between various methods with human judgments. Our MetaCritique Meta-F1 beats the Single baseline by a large margin, confirming its increased reliability. Also, we observe that Meta-P has a reduced correlation in human-written critiques, likely because humans make fewer factuality mistakes, resulting in clustered precision scores. Lastly, Meta-Open distinctly exceeds the performance of the Single baseline, demonstrating that even less advanced LLMs can surpass GPT-4 Single baseline. This indicates that our MetaCritique framework is more effective than simple GPT-4 scoring. Lastly, precision and recall scores complement each other in assessing LLM-generated critiques. Meta-P or Meta-R is slightly inferior to Meta-F1.
4.5 Exp-IV: Agreement Rate for Pairwise Comparison
Setup
In this experiment, we use various methods to determine the superior critique out of two that are generated by LLMs. We calculate the agreement rate of different methods consistent with the gold label, which stems from the gold F1 score.
Results
Figure 3 illustrates the agreement rate of various methods. Our MetaCritique Meta-F1 achieves the best performance. Especially, our MetaCritique exceeds the Single baseline by a considerable margin (44%). It implies that the scores from our MetaCritique are more comparable than Single baseline. Even Meta-Open significantly outperforms GPT-4 powered baselines Single and Pairwise, confirming the effectiveness of our framework. Finally, precision and recall scores serve as complementary metrics for evaluating critiques, because Meta-P and Meta-R are somewhat less effective than Meta-F1.
4.6 Exp V: Better Critique, Better Refinement
Setup
Critique is often applied to enhance the quality of model outputs via refinement. It is intuitive that superior critiques result in better refinements. To confirm this hypothesis, we conduct this experiment. Specifically, we instruct GPT-4 to refine the model outputs via the critique. Detailed instructions with demonstrations are presented in Table 17. Subsequently, we compare the refined outcomes to choose the better one. We conduct GPT-4 evaluation and human evaluation for this comparison. The prompt for GPT-4 evaluation is shown in Table 18, while the equivalent instruction for human evaluators is outlined in Table 19.
Results
We report the win rate of refined outcomes from superior critique over inferior critique, where the critique is evaluated by different methods. The outcomes of GPT-4 evaluation are depicted in Figure 4, while the results of the human evaluation appear in Figure 5. The superior critique chosen by our MetaCritique enhances refinement significantly. Moreover, precision and recall scores are mutually supportive. Relying solely on one metric can lead to diminished performance.
5 MetaCritique Leaderboard
We used MetaCritique to rank the critique models, such as GPT 3.5, SelFee (13B), UltraCM (13B), AUTO-J (13B), and human annotators stemming from Shepherd dataset. Evaluating each model for 100 critiques incurs a cost of approximately $50 when utilizing the GPT-4 API.
The results are shown in Table 6. AUTO-J is the best open-source critique model, delivering more factual and comprehensive feedback than its open-source counterparts. Also, it shows that AUTO-J surpasses other critique models in both MetaCritique recall and F1 score. In contrast, human and GPT-3.5 achieve precision scores exceeding 80%, outshining the performance of open-source critique models. This finding highlights that the research of open-source critique models should pay more attention to factuality issues.
| MetaCritique | |||
| Models | Meta-P | Meta-R | Meta-F1 |
| Human | 83.19 | 60.65 | 64.02 |
| GPT 3.5 | 80.79 | 64.27 | 68.72 |
| SelFee | 69.56 | 51.05 | 54.22 |
| UltraCM | 73.64 | 66.77 | 67.79 |
| AUTO-J | 76.43 | 70.65 | 71.14 |
6 Related Work
6.1 Critique Evaluation
In light of the rapid advancements in LLMs, the significance of generating critique is increasingly acknowledged by researchers (Saunders et al., 2022; Wang et al., 2023b; Gou et al., 2023; Madaan et al., 2023; Kim et al., 2023; Ye et al., 2023; Wang et al., 2023a; Welleck et al., 2023; Li et al., 2023; Ke et al., 2023; Cui et al., 2023). However, the field is hindered by the scarcity of adequate research on critique evaluation, a gap that is currently constraining the evolution of this domain. Critique evaluation aims to evaluate the quality of the critique. It mainly relied on the human annotators (Saunders et al., 2022; Wang et al., 2023a), which entails considerable costs and carries a substantial risk of subjectivity. Wang et al. (2023a) used GPT-4 to replace human annotators, but it still lacks transparency since the numeric score is produced directly via GPT-4 without any fine-grained calculation explanation.
6.2 Meta Evaluation
Meta evaluation is designed to assess automated metrics by determining the degree of correlation between automated scores and human evaluations (Zhang* et al., 2020; Yuan et al., 2021; Sai et al., 2022; Fu et al., 2023). This is achieved by utilizing correlation coefficients. The Spearman correlation (Zar, 2005) evaluates the monotonic relationship between two variables, focusing on their ranked values rather than raw data. Conversely, the Pearson correlation (Lee Rodgers and Nicewander, 1988; Mukaka, 2012) is used to gauge the linear relationship between two variables, employing the actual data values. Furthermore, Kendall’s Tau (Kendall, 1938) is utilized to ascertain the ordinal association between two quantified variables. Besides, the significance test (Williams., 1959; Koehn, 2004) served as a crucial supplementary technique to measure the improved correlations. Lastly, accuracy sometimes serves as meta evaluation to measure the percentage of correct ranking between two outcomes (Yuan et al., 2021).
Meta evaluation is also devised to evaluate the robustness of metrics (Zhang* et al., 2020; Sai et al., 2022). Traditional metrics, which primarily rely on lexical attributes such as word overlap, can be easily attacked by text manipulation. These evaluation methods focused on adversarial scenarios or corner cases in which they fail (Callison-Burch et al., 2006; Hodosh and Hockenmaier, 2016; Kilickaya et al., 2017; Li et al., 2017; Sai et al., 2019). Recently, Zhang* et al. (2020) evaluated the robustness by using the task of adversarial paraphrase classification. The advent of generative AI has greatly mitigated this problem. Thus, we will not pursue this line of research.
6.3 Factuality Detection
Factuality detection aims to classify whether a textual statement, termed claim, is factual (Wang, 2017; Thorne et al., 2018; Augenstein et al., 2019; Zhong et al., 2020; Wadden et al., 2020; Guo et al., 2022). The FEVER dataset, introduced by Thorne et al. (2018), has led to the development of fact-checking models (Zhong et al., 2020; Krishna et al., 2022). In this context, only the claims from Wikipedia are provided to verify the claim without related evidence. Besides, factuality detection can also be formalized as factual consistency in summarization tasks, such as FactCC (Kryscinski et al., 2020) and QAGS-based models (Wang et al., 2020), which determines whether the produced summaries or summary sentences align factually with the provided evidence. In this task setting, the claim and its corresponding evidence are given. Similarly, WICEbased methods (Kamoi et al., 2023) aim to classify whether a given claim from Wikipedia can be entailed by the provided evidence. Lastly, Gao et al. (2023) and Chern et al. (2023) detected factuality without the claim and evidence. RARR (Gao et al., 2023) prompted LLMs to generate claims, retrieve evidence, and verify the claim. FacTool (Chern et al., 2023) utilized LLMs with different tools to detect factuality in various scenarios.
7 Conclusion
We are the pioneers in prioritizing critique evaluation and introducing the critique of critique, termed MetaCritique. We introduce two principles of precision and recall for the critique evaluation. Our MetaCritique is quantitative and interpretable, wherein critiques are decomposed into AIUs, and two pragmatic tasks are established to calculate objective numeric scores along with plausible natural language rationale. We collected a meta-evaluation dataset involving question answering, reasoning, entailment, and summarization to confirm the feasibility and effectiveness of our MetaCritique. Experiments also demonstrate that superior critiques chosen by our MetaCritique can result in a better refined outcome, suggesting that our MetaCritique holds the potential to substantially enhance generative artificial intelligence.
8 Limitation
The creative tasks are not suitable for recall principle, especially when there are multiple high-quality answers. It is intuitive that the suggestions for these tasks are various, so the critiques are also diverse. As a result, it is very hard to collect all effective critiques for recall principle.
Additionally, reference answers or reference critiques are often unavailable. Within our setup, GPT-4 serves as a reference, yet it is important to acknowledge that GPT-4 may commit errors.
In future work, it is essential to expand upon exploring open-source large language models to achieve more robust and dependable evaluations. This also leads to reliability and numerical stability, through the transparent and verifiable process inherent in open-source developments.
Appendix A Prompts and Instructions
We detail the prompts for GPT-4 generation, including generating hypothesis/reference critique, extracting AIUs, performing precision/recasll task, refining model-generated answer, implementing GPT-4 baselines (Single and Pairwise), and evaluating refined answer. Additionally, we provide guidelines for human evaluation, covering generated outcomes such as AIUs, reference answer/critique, refined answer. For creating reference answers, the system prompt employed in our paper is “You are a brilliant AI assistant.” We used “claims” instead of “AIUs” in our prompts to reduce possible ambiguity and confusion.
————–SYSTEM MESSAGE————- You are a brilliant AI assistant. You will receive an input question and the model-generated answer. You need to generate the specific and actionable critiques for the model-generated answer, which contain the critical comments and helpful suggestions. ————–USER MESSAGE————- input question: {question} model-generated answer: {model-generated answer} critique: |
You are a brilliant AI assistant. You will receive an input question and the model-generated answer. You need to generate the specific and actionable critiques for the model-generated answer, which contain the critical comments and helpful suggestions. input question: {question} model-generated answer: {model-generated answer} critique: |
————–SYSTEM MESSAGE————- You are a brilliant AI assistant. You receive a critique as the input, which is the critical comment for an input question and a model-generated answer. You need to generate all check-worthy claims of the critique. A claim is an atomic statement that only contains a fine-grained information unit of a critique. Each claim should be concise (a sentence) and self-contained. Note that the ’answer’ in the critique means the model-generated answer or the correct answer. Some examples are as following: [example 1 start] input: The answer violates basic principles of common sense. Flour is not sweet. Dusting it onto the bread would not sweeten the bread. Therefore, the right answer is to dust powdered sugar; sugar is, of course, sweet. claims: The model-generated answer violates basic principles of common sense that flour is not sweet. Dusting Flour onto the bread would not sweeten the bread. The correct answer is to dust powdered sugar. Sugar is sweet. [example 1 end] [example 2 start] input: The output makes a logical error in the first bullet point of the answer, where it rejects the possibility of sunlight being the right answer. While sunlight might be produced in the sun, it doesn’t stay there. Since sunlight comes to earth, it is on earth itself. Therefore sunlight, option 2, is the right answer as sunlight which is the oldest heat source on Earth. claims: The model-generated answer makes a logical error in rejecting the possibility of sunlight being the right answer. Sunlight might be produced in the sun. Sunlight doesn’t stay in the sun. Sunlight comes to earth. Sunlight is on earth itself. Sunlight, option 2, is the right answer as it is the oldest heat source on Earth. [example 2 end] [example 3 start] input: That the increase has "reached record levels" is not indicated in the context. Otherwise, the answer is adequate, except it should also mention the government is responding to the situation. claims: The content that the increase has "reached record levels" is not indicated in the context. The correct answer should also mention the government is responding to the situation. [example 3 end] [example 4 start] input: The answer’s summary was missing information about Andre Ward winning the rematch. claims: The model-generated answer was missing information about Andre Ward winning the rematch [example 4 end] ————–USER MESSAGE————- input: {critique} claims: |
————–SYSTEM MESSAGE————- You are a brilliant AI assistant. You receive an input question, a model-generated answer, the reference answer and a claim that is the critique for the model-generated answer. To verify the claim, you should think step by step as following: 1) you should consider the input question, the model-generated answer and the reference answer, and then you should determine what you need for verifying the claim. Please note that the input question is the most reliable, and the input question is always correct. The model-generated answer is the least reliable, and it can contain some flaws. The reference answer is more reliable than the model-generated answer, but sometimes it may contain some flaws. 2) verify whether the claim is true or false. If a claim is either a factual statement, explicitly mentioned in the needed information, or logically deducible from it, then the claim is true; otherwise, it is false. Here are some examples: [example 1 start] <input question> A shelf would fit in what part of a room? Please choose the correct option. Here are the options: Option 1: garage Option 2: corner <model-generated answer> Option 1 <reference answer> Option 2 <claim> The model-generated answer is wrong. <verify claim> To verify the claim, we need to consider the model-generated answer and the reference answer. The reference answer says that Option 2 is the correct choice. The model-generated answer states that the answer is Option 1. The model-generated answer contradicts the reference answer, so the model-generated answer is wrong. The claim states that the model-generated answer is wrong. The claim is consistent with the above analysis result. Therefore, the claim is true. [example 1 end] [example 2 start] <input question> Give a summary of the below article: Biggin Hill Airport wants to increase bank holiday and weekend opening times to increase the number of flights. Director William Curtis said he feared the airport could be asked to help the UK meet increased demand for air travel if the extension was not granted. <model-generated answer> Biggin Hill Airport intends to expand its hours to accommodate more flights. <reference answer> Biggin Hill Airport seeks to extend opening times to boost flights amidst concerns of meeting UK air travel demand. <claim> William Curtis wants to increase opening hours. <verify claim> To verify the claim, we need to consider the input question. The input question says that Biggin Hill Airport wants to increase opening times instead of William Curtis. The claim states that William Curtis wants to increase opening hours. The claim contradicts with the above reasoning result. Therefore, the claim is false. [example 2 end] [example 3 start] <input question> Context: He walked in the door and saw her pointing a gun with a laser pointer at his chest. Given the above context, How did he feel? <model-generated answer> He was glad. <reference answer> He was scared for his life. <claim> The correct answer is that he was afraid. <verify claim> To verify the claim, we need to consider the reference answer. The reference answer states that he was scared for his life, which is synonymous with being afraid. The claim states that the correct answer is that he was afraid. The claim is consistent with the above reasoning result. Therefore, the claim is true. [example 3 end] [example 4 start] <input question> Here is a premise: A black and white dog sliding down a slope on a beach. Here is a hypothesis: A dog going uphill to the beach. Does this premise imply the hypothesis? Please justify your answer: <model-generated answer> No, the dog could have been sliding down any slope, not just a sandy beach. <reference answer> No. The premise states that the dog is sliding down a slope on a beach, which indicates a downward movement. The hypothesis suggests that the dog is going uphill to the beach, which indicates an upward movement. These two scenarios are contradictory, so the premise does not support or imply the hypothesis. <claim> The model-generated answer provides a clear and concise explanation. <verify claim> To verify the claim, we need to consider the model-generated answer and the reference answer. The model-generated answer states that the reason is that the dog could have been sliding down any slope. However, it does not mention the direction of the dog’s movement, which is the main point of contradiction. The claim states that the model-generated answer provides a clear and concise explanation. The claim contradicts with the above reasoning result. Therefore, the claim is false. [example 4 end] ————–USER MESSAGE————- <input question> {question} <model-generated answer> {model-generated answer} <reference answer> {reference answer} <claim> {AIU from hypothesis critique} <verify claim> |
————–SYSTEM MESSAGE————- You are a brilliant AI assistant. You receive a reference text and a claim. Note that the ’answer’ in the reference text means the model-generated answer or the correct answer. You need to think step by step for verifying whether the claim is true or false. If a claim is either explicitly mentioned in the reference text or logically deducible from it, then the claim is true; otherwise, it is false. Here are some examples: [example 1 start] <reference text> The model-generated answer gives the wrong step. It is suggested to check the reasoning steps. <claim> The model-generated answer is wrong. <verify claim> The reference text points that the model-generated answer contains the wrong step, so the model-generated answer is wrong. The claim states that the model-generated answer is wrong. It is consistent with the above reasoning result. Therefore, the claim is true. [example 1 end] [example 2 start] <reference text> The model-generated answer is not right, because it provides the wrong analysis. The correct answer should be Option A. <claim> The correct answer should point out the invented fact. <verify claim> The reference text states that the correct answer should be Option A. The claim states that the correct answer should point out the invented fact. The claim is not mentioned or implied in the reference text. Therefore, the claim is false. [example 2 end] [example 3 start] <reference text> The answer lacks detail and context, like the age of the victim, the place of the incident and so on. <claim> The accurate summary should contain the age of the dead person. <verify claim> The reference text states that the model-generated answer lacks the detail, the age of victim. The claim states that the correct summary should contain the age of the dead person. It means that the model-generated answer is not mentioned the age of victim. The claim can be logically deducible from the reference text. Therefore, the claim is true. [example 3 end] [example 4 start] <reference text> The answer could be more concise and focused. <claim> The model-generated answer is mostly correct, but it could be improved by providing more specific details. <verify claim> The reference text states that the model-generated answer could be more concise. It means that the model-generated answer is elaborated. The claim states that the model-generated answer could be improved by providing more specific details. It means that the model-generated answer is brief. The claim contracts with the reference text. Therefore, the claim is false. [example 4 end] ————–USER MESSAGE————- <reference text> {hypothesis critique} <claim> {AIU from reference critique} <verify claim> |
Your task is to evaluate the generated reference answer (i.e., response for a question). Give a score 1-7 (worst-best) based on the quality of the answer. 7: Perfect. The response perfectly answer the query and provide suitable explanation. 6: Exceptional. The response perfectly answer the query and provide suitable explanation, but introduce some flaws. 5: Excellent. The response correctly answer the query and provide suitable explanation, but introduce some flaws. 4:Good. The response correctly answer the query but lacks suitable explanation. 3: Average. The response incorrectly answer the query but provides consistent explanation. 2: Poor. The response incorrectly answer the query but provides the explanation with flaws. 1: Extremely bad. The generated response is random text or simple repeats the question. Give a score 1-3 for response with incorrect content and give a score 4-7 for response with correct content. |
Your task is to evaluate the critique on a model-generated answer. Give a score 1-7 (worst-best) based on the quality of the critique. 7: When the answer is wrong, the critique clearly highlights the most important errors and provides very actionable suggestions. When the answer is correct, the critique confirms the answer is correct and provides very useful suggestions. 6: When the answer is wrong, the critique confirms that the answer is wrong and points out the most important errors. When the answer is correct, the critique confirms the answer is correct and provides useful suggestions. 5: When the answer is wrong, the critique misses the important errors but clearly confirms that the answer is wrong. When the answer is correct, the critique confirms the answer is correct and proposes some less useful suggestions. 4: The critique has a correct judgment of the answer (e.g., states correct answer is correct or states wrong answer is wrong). 3: The critique is vague about whether or not the answer is correct. Or the critique itself tries to answer the question regardless of the content in the answer. 2: The critique has a wrong judgment of the answer (e.g., states correct answer is wrong or states wrong answer is correct). 1: The critique is completely random text or simply repeats the answer. First, please check whether the critique has correct or incorrect judgment (correct judgment means the answer is correct, critique confirms the correctness. Or if the answer is incorrect, the critique confirms the incorrectness.) Give a score 1-3 for critique with incorrect judgment and give a score 4-7 for critique with correct judgment. |
Your task is to evaluate the AIU extraction on a critique. Give a score 1-7 (worst-best) based on the quality of the extraction. 7: Perfect. Generate all salient claims without extra information. Useless or redundant information is removed. 6: Exceptional. Generate all salient claims without extra information but contains very few useless or redundant information. 5: Excellent. Generate all salient claims but introduce little extra information. 4:Good. Generate some salient claims but introduce little extra information. 3: Average. Generate some salient claims but introduce too much extra information. 2: Poor. Remove too many salient claims, or introduce too much extra information. 1: Extremely bad. The generated claims are random text or simple repeats the critique. |
————–SYSTEM MESSAGE————- You are a brilliant AI assistant. You will receive a question, a model-generated answer, and two critiques about this answer. A good critique should point out key errors contained in the answer and provide constructive suggestions. Your task is to evaluate the quality of the critique. Your evaluation should consider factors such as the accuracy, factuality, comprehensiveness, relevance, and conciseness. Begin your evaluation by comparing the two critiques and provide a short explanation. Avoid any position biases and ensure that the order in which the answers were presented does not influence your decision. Do not allow the length of the critiques to influence your evaluation. Do not favor certain names of the answers. Be as objective as possible. After providing your explanation, output your final verdict by strictly following this format: "[[A]]" if the critique A is better, "[[B]]" if the critique B is better, and "[[C]]" for a tie. ————–USER MESSAGE————- <input question> {question} <model-generated answer> {model-generated answer} <critique A> {hypothesis critique from LLM 1} <critique B> {hypothesis critique from LLM 2} |
————–SYSTEM MESSAGE————- You are a brilliant AI assistant. You will receive a question, a model-generated answer, and a critique about this answer. Your task is to evaluate the quality of the critique and give a score. The score is based on the quality of the critique: 7: When the answer is wrong, the critique clearly highlights the most important errors and provides very actionable suggestions. When the answer is correct, the critique confirms the answer is correct and provides very useful suggestions. 6: When the answer is wrong, the critique confirms that the answer is wrong and points out the most important errors. When the answer is correct, the critique confirms the answer is correct and provides useful suggestions. 5: When the answer is wrong, the critique misses the important errors but clearly confirms that the answer is wrong. When the answer is correct, the critique confirms the answer is correct and proposes some less useful suggestions. 4: The critique has a correct judgement of the answer (e.g., states correct answer is correct or states wrong answer is wrong). 3: The critique is vague about whether or not the answer is correct. Or the critique itself tries to answer the question regardless of the content in the answer. 2: The critique has a wrong judgement of the answer (e.g., states correct answer is wrong or states wrong answer is correct). 1: The critique is completely random text or simply repeats the answer. Begin your evaluation by considering the critique and provide a short explanation. First, please check whether the critique has correct or incorrect judgment (correct judgment means the answer is correct, critique confirms the correctness. Or if the answer is incorrect, the critique confirms the incorrectness.) Please note that give a score 1-3 for critique with incorrect judgment and give a score 4-7 for critique with correct judgment. Be as objective as possible. After providing your explanation, output your final score by strictly following this format: "[[score]]". ————–USER MESSAGE————- <input question> {question} <model-generated answer> {model-generated answer} <critique> {hypothesis critique from LLM} |
————–SYSTEM MESSAGE————- You are a brilliant AI assistant. Based on the provided critique, please refine the model-generated answer to address all issues in the critique and ensure that all suggestions from the critique are incorporated to enhance the refined answer. Here are some examples: [example 1 start] <input question> A shelf would fit in what part of a room? Here are the options: Option 1: garage Option 2: corner Please choose the correct option and provide the justification. <model-generated answer> Option 1 <critique> The model-generated answer is wrong and lacks the justification. <refined answer> Option 2. Justification: "Corner" refers to a specific location within a room where two walls intersect, making it the appropriate choice for identifying a part of a room where a shelf could be installed. [example 1 end] [example 2 start] <input question> Context: He walked in the door and saw her pointing a gun with a laser pointer at his chest. Given the above context, How did he feel? <model-generated answer> He was pointed by a gun. <critique> The model-generated answer is a good starting point, but it should mention his feeling. <refined answer> He was scared when he saw a gun. [example 2 end] ————–USER MESSAGE————- <input question> {question} <model-generated answer> {model-generated answer} <critique> {hypothesis critique from LLM} <refined answer> |
————–SYSTEM MESSAGE————- You are a brilliant AI assistant. You can evaluate the quality of the model-generated answer for the input question. Your evaluation should consider factors such as the accuracy, factuality, relevance, and conciseness. Please note that details that do not clearly help to answer the input question are considered redundant. Begin your evaluation by comparing the two answers and provide a short explanation. Avoid any position biases and ensure that the order in which the answers were presented does not influence your decision. Do not allow the length of the answers to influence your evaluation. Do not favor certain names of the answers. Be as objective as possible. After providing your explanation, output your final verdict by strictly following this format: "[[A]]" if the model-generated answer A is better, "[[B]]" if the model-generated answer B is better, and "[[C]]" for a tie. ————–USER MESSAGE————- <input question> {question} <model-generated answer A> {refined answer from the LLM 1 critique} <model-generated answer B> {refined answer from the LLM 2 critique} |
Your task is to pick up the better answer from two given answers. Your evaluation should consider factors such as the accuracy, factuality, relevance, and conciseness. Please note that details that do not clearly help to answer the input question are considered redundant. Do not allow the length of the answers to influence your evaluation. Be as objective as possible. Label your final verdict: "A" if the model-generated answer A is better, "B" if the model-generated answer B is better, and "C" for a tie. |